OpenStudy (anonymous):

Find the axis of the graph of the function. f(x) = 4x^2 + 8x + 7

OpenStudy (anonymous):

Is it asking for when Y will be zero? if so replace f(x) with 0 and solve.

OpenStudy (anonymous):

if you mean axis of symmetry thats x=-b/2a when your equation is in the form f(x)=ax^2+bx+c so x=-8/2(4) x=-8/8 x=-1